OpenStudy (anonymous):

Locate the critical point of the function. f(x,y)=2 − x^(2) − 8x − y^(2)+6y Need all the help i can get i have an exam in 4 hours

OpenStudy (amistre64):

find 4 things: Fx, Fy, Fxx, Fyy, Fxy .... ok, 5 things but only the doubles are needed

OpenStudy (anonymous):

i think i got down to the critical points, i found x to be -4 and y to be 3… then do i plug those in?

OpenStudy (amistre64):

D = Fxx Fyy - (Fxy)^2 if D > 0 and Fxx >0, then (a,b) is a relative min if D > 0 and Fxx <0, then (a,b) is a relative max if D < 0, its a saddle point if D = 0, then it may or maynot be sufficient

OpenStudy (amistre64):

let me check .. f(x,y)=2 − x^(2) − 8x − y^(2)+6y Fx = -2x-8 Fxx = -2 Fxy = 0 Fy = -2y+6 Fyy = -2 D = 4 D > 0, Fxx < 0 so there is a max located someplace

OpenStudy (amistre64):

x=4, y=3 seems to fit the bill

OpenStudy (amistre64):

hmm, max = -4,3 yes

OpenStudy (amistre64):

lol, i was thinking -2x+8 .... silly me

OpenStudy (anonymous):

haha its okay, so is that the only two critical points?

OpenStudy (amistre64):

that is 1 critical point, and we found it since this is an upside down parabolic cone

OpenStudy (anonymous):

how did you find out it was an upside down parabolic cone? cause i need to know how to graph on my exam